answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

If you have 15 cards with different shapes If the first card you draw is a square and the probability is 1 in 5 What is the probability of next card something other than a square?

1 in 5 is a square so there are 3 squares and 12 non-squares. If the square is not replaced, and the next card is drawn at random the probability of a non-square is 12/14.

A fair die is rolled 9 times compute the probability that a one appears exactly once?

The probability of a one being rolled in a fair die is 1 in 6, or 0.1666... . The probability of a one not being rolled is 5 in 6, or 0.8333... . The probability, then, of exactly one one being rolled in nine rolls is 1 in 6 times 5 in 6 to the 8th power, or about 0.0388.

A man takes a step forward with probability 0.4 and backwards with probability 0.6 Find the probability that at the end of eleven steps he is just one step away from the starting point?

1. He takes 6 steps forward and 5 steps backward

2. He takes 5 steps forward and 6 steps backward

consider 1:

p1 = (11C6)(0.4)^6*(0.6)^5

for 2:

p2 = 11C5(0.4)^5*(0.6)^6

total probability = p1 + p2

When a single card is drawn from a standard 52 card deck find the probability that the card is a club or a seven?

There are 13 clubs (including the seven of clubs), and 3 other sevens (from the other suits). So that makes 16 cards that are clubs or sevens. Thus, the probability of drawing a club or a seven from a standard 52 card deck is 16/52 = 4/13 (about 31%).

What is the probability of rolling a 4 or an even number on a number cube?

That depends on how many sides the cube has but if it was a 6 sided cube then the probability of getting a four is 1 to 6 and even is 3 to 6.

Since 4 is even, then the result set of 4 intersect with the result set of even, so the probability of rolling a 4 or an even number is 1 in 2, or 0.5.

What happens if your v5 Tamagotchi never married?

You probably have to reset.. I don't know why you woudntmarry though

What percentage chance to you have at being murdered?

The answer depends, among other factors, on what your occupation is and where in the world you live. For most people, the probability is negligible.

What is the theoretical probability of rolling two dice 100 times?

The theoretical probability is 1. Lots of people have done it in the course of their lives so it is an event that has happened and will happen again.

What are the deck odds for a pack of playing cards?

There are twenty odd cards in a deck of cards. This is a math problem.

What are your chances of conceiving twins after you miscarried twins?

I was told by my doctor, that the chances of having multiples is very likely once you've conceived multiples in the past. And if you have multiples, they will skip a generation and then their kids will likely have multiples. Pretty cool.

What occurred in 472 bc?

Ancient Olympic Games were extended from a one day festival with a few events to a five day event with wrestling, boxing, horse racing, jumping, javelin, and chariot racing.

A fair die is rolled 4 times in succession what is the probability that even numbers are rolled all 4?

You want: P(even) * P(even) * P(even) * P(even). P(even) = 3/6 or 1/2. Therefore probability even on 4 rolls is: 1/2 * 1/2 * 1/2 * 1/2 = 1/16 = 0.0625.

What is the probability of getting 2 prime nos from 1 to 20?

In the sample space [1,20], there are 8 prime numbers, [2,3,5,7,11,13,17,19]. The probability, then, of randomly choosing a prime number in the sample space [1,20] is (8 in 20), or (2 in 5), or 0.4. The probability of choosing two of them is (8 in 20) times (7 in 19) which is (56 in 1064) or (7 in 133) or about 0.05263.

Trending Questions
Is drawing a card from a deck replacing it and drawing a second card an example of an Independent Event? Both number cubes with numbers 1 through 6 are rolled what is the probability the white cube will land on 4 and the orange cube also will land on 4? Rolling three dice what is the probability of obtaining at least one six? The probability of rolling an even number when number cube is rolled twice? A committee consists of 7 women and 4men if a member of the committee is chosen at random to act as chairperson what is the probability that the choices is woman? How many different four-letter initials are there that begin with A? Suppose a random variable Z takes on the values 1 2 3 4 and 5 what is the expected value of Z defined as? What are the odds of rolling a seven with two dice fifty six times in succession? How many ways can letters of thousands arranged? How many ace of spade are in a deck of 52 cards? Where do you go to get help answering Statisctic problems? What is 70 out of 200 simplified? What is the volume of 4 marbles using water displacement? What are the events that occur at a neuromuscular junction? How many red card are in a deck of 52? How many ways can you arrange four letters? What is the probability of randomly choosing a vowel from the letters a b c d e? How many ways can a president and vice president be randomly selected from a class of 28 students? What are the chances of a 14 year old boy impregnating a 14 year old girl? What is the probability of getting a sum of 7 or 11 by rolling two dice if the first roll is a 3?